26dec05 A1312 no power polB for 1 file
On 26dec05 project a1312
took two files of OH data with lbw, circular polarization, and the 1550-1820
Mhz filter in. The first file had no power in the correlator polB
data. The second file was ok.
The plot shows the
status of the data during the first file (pdf) (corfile.26dec05.a1312.1):
Top: Lag 0 total power versus time. Black is polA, red is polB. The vertical
scale is linear in power. The value 1 has the measured power equal to the
optimum power. PolB (red) has no power.
2nd: The correlator attenuator values versus time. The polB attenuators
where at the minimum value (0) since the input power was low.
3rd: The power measured at the input to the upstairs fiber optic transmitters.
PolA and polB are similar. The value should be close to -40 dbm (which
4th: The power at the top of the downstairs iflo (after the fiber optic
receiver and the variable amps. This value was only read once (at the very
start of the run). Both values were reasonable. We don't know whether this
changed during the run.
The dewar/first lo is probably ok since the upstairs power meter read ok
for both pols. It is possible that the incorrect filter was inserted in
polB. The if1 power meter reads about 500 Mhz bandwidth so it is possible
that it read ok and then the correlator 3 Mhz bandpass was out of band.
This is doubtful since the upstairs attenuation for polA and polB were
The correlator power counters (50 MHz bandwidth) also had no power. The
power counters are analog and they are before the correlator digitizers.
So this is not a problem with the half band filters and probably not the
The downstairs iflo power meter was only read once at the very start of
the run (when the command adjpwrif2 was executed). At the beginning the
input power to the downstairs if/lo was ok. This could have changed
later in the observation.
It was probably not a problem with the downstairs synthesizers since each
synth supplies power to both pol A and polB simultaneously.
My guess is that one of the downstairs switches on the polB side did not
setup correctly causing no power to pass to the correlator. When the later
observation was done, the iflo was resetup and things worked correctly
(although this is only a guess).